OpenStudy (anonymous):

the number of group isomorphic from s3 to z/6z

OpenStudy (kinggeorge):

To restate, you're asking for the number of isomorphisms \(S_3 \longrightarrow \mathbb{Z}/6\mathbb{Z}\) Where \(S_3\) is the permutation group of 3 elements correct?

OpenStudy (anonymous):

yes

OpenStudy (kinggeorge):

Let's think about this a little. What's one of the key properties that \(\mathbb{Z}/6\mathbb{Z}\) has?

OpenStudy (anonymous):

is it a cyclic group z/nz is isomorphic to nz

OpenStudy (kinggeorge):

Both important properties, but in this case we know that it's also abelian.

OpenStudy (kinggeorge):

As for \(S_3\), it is not abelian. Thus, the two groups are not isomorphic.

OpenStudy (anonymous):

that means no of isomorphisms is zero

OpenStudy (kinggeorge):

yes.

OpenStudy (anonymous):

thanks will think about it bye